# When Should Selfish Miners Double-Spend?

### v3

@@ 5 Analysis of $S$-Stealth Mining
An interesting observation of the results in Theorem 3 is the fact that $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that, even if the attack cycles end successfully, $\gamma$ increases the probability that the prefix of $H$-chain switches to $tx^{\prime}$ before $tx$ is confirmed, because the adversary matches each honest block before $A=L$. Thus, for a given $\alpha$, while $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease and $P(\textbf{Move-Funds})$ increases. Further, even if the $H$-chain still contains $tx$ in its prefix, every time an honest block is mined, the adversary matches it, i.e., releases a competing block on the same height which contains $tx^{\prime}$ in its prefix, which would make the merchant providing services $tx$ suspicious. In this section, we provide a modified version of $L$-stubborn mining, which we call $S$-stealth mining, aiming to make the detection of the double-spending attack difficult and to increase $P(\textbf{Double-Spending})$ in each attack cycle.
@@ 5.1 Optimizing the Revenue Ratio
Here, we first provide the theorem stating the quasiconcavity of $S$-stealth mining. Even though the goal in $S$-stealth mining is to increase the double-spending chance instead of focusing on the revenue ratio, for the sake of completeness, for a given $(\alpha,\gamma)$ pair, we also provide an algorithm to find $S^{*}$, i.e., the integer value of $S$ maximizing the revenue ratio. The proofs are provided in Appendix C.
@@ 5.2 $(k+1)$-Stealth Mining and Double-Spending
Proof:  Contrary to $(k+1)$-stubborn mining, a successful cycle of $(k+1)$-stealth mining always results in double-spending since the $H$-chain always keeps $tx$ in its prefix and the $A$-chain undoes it eventually. On the other hand, if an attack ends unsuccessfully with $H=(k+1)$, the honest chain might switch from $tx$ to $tx^{\prime}$ right at the end when the honest block at height $k+1$ is mined. This would imply that the first $tx$ is confirmed when the $H$-chain reaches length $k$ and the next honest block switches prefix and $tx^{\prime}$ is confirmed, resulting in a double-spend.  $\blacksquare$
@@ 5.3 Combined Revenue Ratio
(Combined Revenue) The combined revenue ratio of $(k+1)$-stealth mining and the associated double-spending is given as
@@
Table 1: Comparison between the revenue ratio of $L$-Stubborn mining and MDP models [9, 5] for $\gamma=0$.
@@ 6 Numerical Results
In Table 1, we compare the optimal revenue ratio obtained from $L^{*}$-stubborn mining with the revenue ratio obtained from selfish mining [4] (where $L=2$) and MDP models of Optimal Selfish Mining (OSM)[5] and Probabilistic Termination Optimization (PTO) [9] to demonstrate that our model performs close to the $\epsilon$-optimal selfish mining strategies. We use the data from [9, Table 1], where $\gamma=0$ and both MDP models are limited to maximum fork length $95$. It is easy to verify that our $L$-stubborn strategy behaves close to the optimal solution of MDP models for $\gamma=0$. The small difference can be attributed to the fact that, in MDP models, when the adversarial chain falls behind honest chain while having a length close to but smaller than $L$, adversary does not immediately adopt the honest chain but instead tries to catch up the honest chain which brings some additional revenue, albeit rather small. In Table 2, we compare $\rho_{L^{*}}$ with $\rho_{OSM}$ of [5] for varying $\alpha$ and $\gamma$ values, where we use the implementation of [52] for the MDP of [5] with a maximum fork length of $80$. It is quite clear that our simplified $L^{*}$-stubborn strategy results in reward ratio $\rho_{L^{*}}$, that is usually within $10^{-2}$ of the reward ratio $\rho_{OSM}$.
@@
Similar to the results of $L$-stubborn mining, Figure 2 also contains related results for $S$-stealth mining. Observing the $S^{*}$ provided in Figure 2(b), it is clear that $S^{*}\leq L^{*}$ which is straightforward from the analysis provided earlier, but also $S^{*}\leq 4$ for $\alpha<0.5$ and $\gamma<1$. In other words, in stealth mining, adversary cannot act in a stubborn manner forever, whereas acting stubborn no matter what the length of the adversarial branch is the best strategy to maximize the revenue ratio for $L$-stubborn mining in the white region of Figure 2(a). In Figure 2(d), we provide the normalized revenue ratio of optimal $S^{*}$-stealth mining attack. Although it starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$ similar to the $L^{*}$-stubborn mining, the rise in the revenue ratio is not as steep initially. Further, as expected, the revenue difference between $L^{*}$-stubborn mining and $S^{*}$-stealth mining increases with $\gamma$, as matching is only present in the last step in stealth mining. Finally, Figure 2(f) shows the values of $\bar{S}$, where we have $\bar{S}<\infty$ for all values of $\alpha<0.5$ and $\gamma<1$, since $\sigma_{\infty}=0$.
In Figure 3 we provide the probabilities of the events Double-Spending, Move-Funds and Service for each attack cycle of $(k+1)$-stubborn mining where we pick $k=6$ as in Bitcoin. We also provide normalized revenue ratio of $(k+1)$-stubborn mining in Figure 3(d) where the red curve highlights the boundary where the ratio rises above 1. It is clear that the red curve is the same as the boundary between $\bar{L}=6$ and $\bar{L}=7$ provided in Figure 2(e) by definition. For all values above and right of the red curve, a transaction is at risk of getting reversed in the sense that it is either not going to be confirmed at all or worse, it will be confirmed and then replaced with a conflicting transaction at no-cost to an adversary. The associated risks are provided in Figure 3(b) and Figure 3(a), respectively. As it is clear from Theorem 3, $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that even if the attack cycles end successfully, $\gamma$ increases the probability of Move-Funds event as the prefix of the $H$-chain switches to $tx^{\prime}$ before $tx$ is confirmed because the adversary matches each honest block before $A=L$. Hence, for a given $\alpha$ as $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease, whereas $P(\textbf{Move-Funds})$ increases. Further, as mentioned earlier, $P(\textbf{Service})>\beta$ since the first honest block contains $tx$ and becomes part of the offset chain as soon as it is mined with probability $\beta$. If one were to consider pre-mining strategies, $P(\textbf{Double-Spending})$ could be increased at the expense of $P(\textbf{Service})$, hence security threats can be even more imminent than we display here.
@@ 7.1 Discussion
Notice that, in the literature, the revenue loss of the adversary is usually ignored when considering the double-spending attacks except for the MDP models of [8, 7] where a stochastic explicit analysis is missing and MDPs are used as black boxes to solve the problem. On the other hand, studies such as [16] focus on the double-spend value and only consider the usual double-spending attack structure of [3]. The attack cycles we describe and analyze in this paper allow an attacker to mount a sustained double-spending attack until it succeeds while protecting the attacker from revenue loss in terms of the coinbase rewards in the long run by combining the attack with selfish mining. Here, it is worth noting the severity of the selfish mining and double-spending attacks in general. The selfish mining threatens the fairness of the incentives. A selfish mining attack decreases the revenue and revenue ratio of all honest miners. Under the attack, if the honest mining is not sustainable anymore when the mining costs are taken into account, honest miners leave the system, increasing attackers fraction of hashrate and destabilizing the entire system. Even if the honest mining is sustainable despite the unfair reward distribution, the attacker pool grows by rational miners joining it, destabilizing the entire system again.
@@ 7.2 Extensions
There are two possible extensions to our paper that can further increase the action space and combining the two would eventually allow all possible state-space considered in [5]. First, note that, we define the attack cycles in a way such that they are i.i.d. This is a design choice which allows us to know whether $tx$ is confirmed or not and use Doob’s optional stopping theorem in our analysis as well as the techniques provided in Appendix B and Appendix C. In this paper, each time an attack cycle starts, we have $A=0$ and $H=0$, since we essentially keep track of the length of the $A$-chain, even if the $H$-chain has some prefix consisting of adversarial blocks of the $A$-chain. On the other hand, one could consider the situations where the duration of attack cycles are defined with respect to the change in the prefix of the $H$-chain, i.e., restricting $H$-chain to consist of the honest blocks exclusively, which is essentially considered in the MDP defined in [5]. This change not only complicates the rigorous analysis and requires us to keep a parameter that tracks the number of honest blocks, but also requires us to keep track of the depth of the $tx$ and $tx^{\prime}$, hence our design choice is to avoid this. The other extension is more obvious, which is to allow trail stubbornness, i.e., to allow the adversary to continue the attack even if it falls behind. By considering an extra parameter for the upper bound on the deficit of the $A$-chain compared to the $H$-chain, our analysis can be redone, however, this would also complicate the analysis provided in Appendix B and Appendix C. Note that, increasing the action space for the attacker increases the maximum revenue ratio that can be obtained. However, based on our simulations and comparison with the maximum revenue ratios obtained by the MDP of [5] for several $(\alpha,\gamma)$ parameters, the difference is negligible.
Appendix A Selfish Mining [4] and OSM MDP [5]
Here, we restate and summarize the important bits of the selfish mining attack of [4] as well as the Optimal Selfish Mining (OSM) MDP model introduced in [5] to find the $\epsilon$-optimal selfish mining strategies.
A.1 Selfish Mining of Eyal and Sirer [4]
@@
    If the attacker pool mines a block first, it hides it and tries to extend it. When the next block arrives:
@@
      If the block is adversarial, the attacker has $2$ private blocks. It will keep mining privately until the difference between the public branch and private branch reduces to $1$, at which point it will release the private branch to override the public branch.
@@ A.2 OSM MDP of [5]
$\rho$ $=\mathbb{E}\left[\liminf_{T\rightarrow\infty}\frac{\sum_{t=1}^{T}r^{t}_{1}}{{\sum_{t=1}^{T}r^{t}_{1}+\sum_{t=1}^{T}r^{t}_{2}}}\right].$ (43)
@@
$v^{\pi}_{\rho}$ $=\mathbb{E}\left[\liminf_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}w_{\rho}(r^{t}_{1},r^{t}_{2})\right].$ (45)
@@
Appendix B Proof of Theorem 2 and Algorithm 1

### v4

@@ 5 Analysis of $S$-Stealth Mining
An interesting observation of the results in Theorem 3 is the fact that $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that, even if the attack cycles end successfully, $\gamma$ increases the probability that the prefix of $H$-chain switches to $\mathsf{B^{\prime}}$ before $\mathsf{B}$ is confirmed, because the adversary matches each honest block before $A=L$. Thus, for a given $\alpha$, while $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease and $P(\textbf{Move-Funds})$ increases. Further, even if the $H$-chain still contains $\mathsf{B}$ in its prefix, every time an honest block is mined, the adversary matches it, i.e., releases a competing block on the same height which contains $\mathsf{B^{\prime}}$ in its prefix, which would make the merchant providing services related to $\mathsf{B}$ suspicious. In this section, we provide a modified version of $L$-stubborn mining, which we call $S$-stealth mining, aiming to make the detection of the double-spending attack difficult and to increase $P(\textbf{Double-Spending})$ in each attack cycle.
@@ 5.1 Optimizing the Revenue Ratio
Here, we first provide the theorem stating the quasiconcavity of $S$-stealth mining. Even though the goal in $S$-stealth mining is to increase the double-spending chance instead of focusing on the revenue ratio, for the sake of completeness, for a given $(\alpha,\gamma)$ pair, we also provide an algorithm to find $S^{*}$, i.e., the integer value of $S$ maximizing the revenue ratio. The proofs are provided in Appendix D.
@@ 5.2 $(k+1)$-Stealth Mining and Double-Spending
Proof:  Contrary to $(k+1)$-stubborn mining, a successful cycle of $(k+1)$-stealth mining always results in double-spending since the $H$-chain always keeps $\mathsf{B}$ in its prefix and the $A$-chain undoes it eventually. On the other hand, if an attack ends unsuccessfully with $H=(k+1)$, the honest chain might switch from $\mathsf{B}$ to $\mathsf{B^{\prime}}$ right at the end when the honest block at height $k+1$ is mined. This would imply that at first $\mathsf{B}$ is confirmed when the $H$-chain reaches length $k$ and the next honest block switches prefix and $\mathsf{B^{\prime}}$ is confirmed, resulting in a double-spend. $\blacksquare$
@@ 5.3 Combined Revenue Ratio
(Combined Revenue Ratio) The combined revenue ratio of $(k+1)$-stealth mining and the associated double-spending is given as
@@
Table 1: Comparison between the revenue ratio of $L^{*}$-stubborn mining and MDP models [9, 5] for $\gamma=0$.
@@ 6 Numerical Results
In Table 1, we compare the optimal revenue ratio obtained from $L^{*}$-stubborn mining with the revenue ratio obtained from selfish mining [4] (where $L=2$) and MDP models of Optimal Selfish Mining (OSM)[5] and Probabilistic Termination Optimization (PTO) [9] to demonstrate that our model performs close to the $\epsilon$-optimal selfish mining strategies. We use the data from [9, Table 1], where $\gamma=0$ and both MDP models are limited to maximum fork length of $95$. In Table 2, we compare $\rho_{L^{*}}$ with $\rho_{OSM}$ of [5] for varying $\alpha$ and $\gamma\neq 0$ values, where we use the implementation of [53] for the MDP of [5] with a maximum fork length of $80$. It is quite clear that our simplified $L^{*}$-stubborn strategy results in reward ratio $\rho_{L^{*}}$, that is usually within $10^{-2}$ of the reward ratio $\rho_{OSM}$ for varying values of $\gamma$. As we mentioned earlier, the small difference can be attributed to the fact that, in MDP models, when the adversarial chain falls behind honest chain while having a length close to but smaller than $L$, the adversary does not immediately adopt the honest chain but instead tries to catch up to the honest chain which brings some additional revenue, albeit rather small.
@@
Similar to the results of $L$-stubborn mining, Figure 2 also contains related results for $S$-stealth mining. Observing the $S^{*}$ provided in Figure 2(b), it is clear that $S^{*}\leq L^{*}$ which is straightforward from the analysis provided earlier, but also $S^{*}\leq 4$ for $\alpha<0.5$ and $\gamma<1$. In other words, in stealth mining, the adversary cannot act in a stubborn manner forever, whereas acting stubborn no matter what the length of the adversarial branch is the best strategy to maximize the revenue ratio for $L$-stubborn mining in the white region of Figure 2(a). In Figure 2(d), we provide the normalized revenue ratio of optimal $S^{*}$-stealth mining attack. Although it starts as $1$ from $\alpha=0$ and reaches $2$ at $\alpha=0.5$ similar to the $L^{*}$-stubborn mining, the rise in the revenue ratio is not as steep initially. Further, as expected, the revenue difference between $L^{*}$-stubborn mining and $S^{*}$-stealth mining increases with $\gamma$, as matching is only present in the last step in stealth mining. Finally, Figure 2(f) shows the values of $\bar{S}$, where we have $\bar{S}<\infty$ for all values of $\alpha<0.5$ and $\gamma<1$, since $\sigma_{\infty}=0$.
In Figure 3 we provide the probabilities of the events Double-Spending, Move-Funds and Service for each attack cycle of $(k+1)$-stubborn mining where we pick $k=6$ as in Bitcoin. We also provide normalized revenue ratio of $(k+1)$-stubborn mining in Figure 3(d) where the red curve highlights the boundary where the ratio rises above 1. It is clear that the red curve is the same as the boundary between $\bar{L}=6$ and $\bar{L}=7$ provided in Figure 2(e) by definition. For all values above and right of the red curve, a transaction is at risk of getting reversed in the sense that it is either not going to be confirmed at all or worse, it will be confirmed and then replaced with a conflicting transaction at no-cost to an adversary. The associated risks are provided in Figure 3(b) and Figure 3(a), respectively. As it is clear from Theorem 3, $P(\textbf{Double-Spending})$ is decreasing in $\gamma$ due to the fact that even if the attack cycles end successfully, $\gamma$ increases the probability of Move-Funds event as the prefix of the $H$-chain switches to $\mathsf{B^{\prime}}$ before $\mathsf{B}$ is confirmed because the adversary matches each honest block before $A=L$. Hence, for a given $\alpha$ as $\gamma$ increases, both $P(\textbf{Double-Spending})$ and $P(\textbf{Service})$ decrease, whereas $P(\textbf{Move-Funds})$ increases. Further, as mentioned earlier, $P(\textbf{Service})>\beta$ since the first honest block $\mathsf{B}$ becomes part of the offset chain as soon as it is mined with probability $\beta$. If one were to consider pre-mining strategies, $P(\textbf{Double-Spending})$ could be increased at the expense of $P(\textbf{Service})$, hence security threats can be even more imminent than we display here.
@@ 7.1 Discussion
Notice that, in the literature, the revenue loss of the adversary is usually ignored when considering the double-spending attacks except for the MDP models of [8, 7] where a stochastic explicit analysis is missing and MDPs are used as black boxes to solve the problem. On the other hand, studies such as [16] focus on the double-spend value and only consider the usual double-spending attack structure of [3]. The attack cycles we describe and analyze in this paper allow an attacker to mount a sustained double-spending attack until it succeeds while protecting the attacker from revenue loss in terms of the coinbase rewards in the long run by combining the attack with selfish mining. Here, it is worth noting the severity of the selfish mining and double-spending attacks in general. The selfish mining threatens the fairness of the incentives. A selfish mining attack decreases the revenue and revenue ratio of all honest miners. Under the attack, if the honest mining is not sustainable anymore when the mining costs are taken into account, honest miners leave the system, increasing attackers fraction of hashrate and destabilizing the entire system. Even if the honest mining is sustainable despite the unfair reward distribution, the rational miners can profit more if they join the adversary, destabilizing the entire system again.
@@ 7.2 Extensions
There are two possible extensions to our paper that can further increase the action space and combining the two would eventually allow all possible state-space considered in [5]. First, note that, we define the attack cycles in a way such that they are i.i.d. This is a design choice which allows us to know whether $\mathsf{B}$ is confirmed or not and use Doob’s optional stopping theorem in our analysis as well as the techniques provided in Appendix C and Appendix D. In this paper, each time an attack cycle starts, we have $A=0$ and $H=0$, since we essentially keep track of the length of the $A$-chain, even if the $H$-chain has some prefix consisting of adversarial blocks of the $A$-chain. On the other hand, one could consider the situations where the duration of attack cycles are defined with respect to the change in the prefix of the $H$-chain, i.e., restricting $H$-chain to consist of the honest blocks exclusively, which is essentially considered in the MDP defined in [5]. This change not only complicates the rigorous analysis and requires us to keep a parameter that tracks the number of honest blocks, but also requires us to keep track of the depth of the $\mathsf{B}$ and $\mathsf{B^{\prime}}$, hence our design choice is to avoid this. The other extension is more obvious, which is to allow trail stubbornness, i.e., to allow the adversary to continue the attack even if it falls behind. By considering an extra parameter for the upper bound on the deficit of the $A$-chain compared to the $H$-chain, our analysis can be redone, however, this would also complicate the analysis provided in Appendix C and Appendix D. Note that, increasing the action space for the attacker increases the maximum revenue ratio that can be obtained. However, based on our simulations and comparison with the maximum revenue ratios obtained by the MDP of [5] for several $(\alpha,\gamma)$ parameters, the difference is negligible.
Appendix A Stochastic Tools for Analysis
A.1 Catalan Numbers and Dyck Words
$C_{n}=\frac{1}{n+1}{2n\choose n}$. (38)
Catalan numbers also satisfy the following recursion
$C_{n}=\frac{2(2n-1)}{n+1}C_{n-1}$ (39)
with $C_{0}=1$. The generating function of the Catalan numbers is
$C(x)=\sum_{n\geq 0}C_{n}x^{n}=\frac{2}{1+\sqrt{1-4x}}.$ (40)
Pre-Dyck words satisfy the recursion
$P[n,m]=P[n,m-1]+P[n-1,m].$ (41)
Clearly $P[n,m]=0$ if $m>n$ and it is easy to verify that $P[n,0]=1$ and $P[n,1]=n$. Thus, (41) gives a straightforward formula for finding $P[n,m]$ in $\mathcal{O}(nm)$.
A.2 Bailey’s Number and Bertrand’s Ballot Problem
$P[n,m]=\frac{n-m+1}{n+m+1}{n+m+1\choose n+1}$. (42)
The elements of the matrix $P$ constitute Catalan’s triangle. Note that, if a broad range of $P[n,m]$ is needed, it is better to start from $P[1,0]=1$ and fill the matrix $P$ via (41). The elements of $P$ matrix appear in [34] where Bailey investigates the sequences that contain $n$ 11’s and $m$ 1-1’s such that the running sum is always non-negative. Bailey proved that the number of such sequences is equal to $P[n,m]$ provided in (42).
A.3 Wald’s Identity and Doob’s Optional Stopping Time
Appendix B Selfish Mining [4], OSM MDP [5] and Our Model
Here, we restate and summarize the important bits of the selfish mining attack of [4] as well as the Optimal Selfish Mining (OSM) MDP model introduced in [5] to find the $\epsilon$-optimal selfish mining strategies. Then, we represent the state and action spaces for $L$-stubborn mining and $S$-stealth mining, which serve as formal descriptions of the strategies introduced in Section 3.1 and Section 3.3 for the sake of completeness.
B.1 Selfish Mining of Eyal and Sirer [4]
@@
    If the adversary mines a block first, it hides it and tries to extend it. When the next block arrives:
@@
      If the block is adversarial, the adversary has $2$ private blocks. It will keep mining privately until the difference between the public branch and private branch reduces to $1$, at which point it will release the private branch to override the public branch.
@@ B.2 OSM MDP of [5]
$\rho$ $=\mathbb{E}\left[\lim_{T\to\infty}\frac{\sum_{t=1}^{T}r^{t}_{1}}{{\sum_{t=1}^{T}r^{t}_{1}+\sum_{t=1}^{T}r^{t}_{2}}}\right].$ (43)
@@
$v^{\pi}_{\rho}$ $=\mathbb{E}\left[\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}w_{\rho}(r^{t}_{1},r^{t}_{2})\right].$ (45)
@@
[…]
B.3.1 $L$-Stubborn Mining
To describe the state space for $L$-stubborn mining, we assume a common block agreed by honest and adversarial miners (either adopt or override action is taken most recently), and the adversary starts a new attack cycle. Then, the state space for each $L$-stubborn mining attack cycle can be succinctly expressed as $(A,H,C,fork)$ where, starting from the common block (excluding), $A$ is the length of the adversarial branch and $H$ is the length of the honest branch and $C$ is the length of the common prefix between the adversarial and honest branches. The field forkfork serves the same purpose as in OSM MDP, i.e., if the last mined block is honest, $fork=relevant$, which implies that the adversary can take the match action whereas $fork=active$ means that a match action is taken recently and $\gamma$ fraction of honest miners are mining on a prefix of adversarial branch of the same length as the honest branch. Under this model, in Table 4, we summarize the state action space and the transitions of $L$-stubborn mining.
State $\times$ Action State Probability Reward
$(A,H,C,\cdot),\text{adopt}^{\dagger}$ $(1,0,0,\text{irrelevant})$ $\alpha$ (C,HC)(C,H-C)
$(0,1,0,\text{irrelevant})$ $1-\alpha$ (C,HC)(C,H-C)
$(A,H,C,\cdot),\text{override}^{\ddagger}$ $(1,0,0,\text{irrelevant})$ $\alpha$ $(A,0)$
$(0,1,0,\text{irrelevant})$ $1-\alpha$ $(A,0)$
$(A,H,C,\text{irrelevant}),\text{wait}$ $(A+1,H,C,\text{irrelevant})$ $\alpha$ $(0,0)$
$(A,H+1,C,\text{relevant})$ $1-\alpha$ $(0,0)$
$(A,H,C,\text{active}),\text{wait}$ $(A+1,H,C,\text{active})$ $\alpha$ $(0,0)$
$(A,H,C,\text{relevant}),\text{match}^{\star}$ $(A,H+1,H,\text{relevant})$ $\gamma\cdot(1-\alpha)$ $(0,0)$
$(A,H+1,C,\text{relevant})$ $(1-\gamma)\cdot(1-\alpha)$ $(0,0)$
$\dagger$ happens whenever $A+1=H$
$\ddagger$ happens whenever $A=H+1\geq L$
$\star$ happens whenever $L>A\geq H$
Table 4: State, action, and transitions of $L$-Stubborn Mining
Notice, in $L$-stubborn mining each state allows a unique action to be taken as presented in Table 4 as opposed to OSM MDP where the goal is to find the optimal action among multiple options for each state. For example, here, whenever $A+1=H$, the action adopt is executed, i.e., no-trail mining is allowed whereas override action is taken whenever $A=H+1\geq L$. Similarly, whenever $fork=relevant$ and $L>A\geq H$, the adversary matches the recently mined honest block. For all the rest of the cases, wait action is taken. Our focus in the analysis is to express the final revenue ratio in terms of $L$ as well as to maximize the revenue ratio with respect to $L$ which essentially changes when each action is taken. Although changing $L$ essentially gives multiple strategies to pick from, this does not cover all the possible strategies as in OSM MDP, hence the final optimal $\rho_{L^{*}}$ is still below $\rho_{OSM}$.
An important observation from Table 4 is that, the value $C$ is only affected by the value of $H$ in a Markovian manner as represented in the last two rows of Table 4. Further, none of the actions depend on $C$ field in the state space and $C$ only affects the reward if an adopt action is taken. This implies that during an $L$-stubborn attack cycle, given $H$, we have
$C=\begin{cases}H-1-i,&\text{with probability }(1-\gamma)^{i}\gamma,\quad 0\leq i<H-1,\\ 0,&\text{with probability }(1-\gamma)^{H-1}.\end{cases}$ (46)
This explains why we treat $C$ as a separate state in the analysis throughout the paper and only introduce here for the sake of completeness. Similarly, since every honest block that is mined on a height where the adversary already has a private block, is matched in $L$-stubborn mining, we did not need to introduce forkfork field in the main body of our paper.
B.3.2 $S$-Stealth Mining
In $S$-stealth mining, match action only happens when in the previous state S1=A=H+1S-1=A=H+1 and a new honest block is mined such that the current state is S1=A=HS-1=A=H at which point the adversary releases its $A$-chain to compete with the honest branch of the same length. As a result, we do not need $C$ and forkfork field in this model and the state $\times$ action space is given in Table 5.
State $\times$ Action State Probability Reward
$(A,H),\text{adopt}^{\dagger}$ $(1,0)$ $\alpha$ $(0,H)$
$(0,1)$ $1-\alpha$ $(0,H)$
$(A,H),\text{override}^{\ddagger}$ $(1,0)$ $\alpha$ $(A,0)$
$(0,1)$ $1-\alpha$ $(A,0)$
$(A,H),\text{wait}$ $(A+1,H)$ $\alpha$ $(0,0)$
$(A,H+1)$ $1-\alpha$ $(0,0)$
$(A,H),\text{match}^{\star}$ $(0,0)$ $\alpha$ $(A+1,0)$
$(0,0)$ $\gamma\cdot(1-\alpha)$ $(A,1)$
$(0,0)$ $(1-\gamma)\cdot(1-\alpha)$ $(0,H+1)$
$\dagger$ happens whenever $A+1=H$
$\ddagger$ happens whenever $A=H+1\geq S$
$\star$ happens whenever S1=A=HS-1=A=H
Table 5: State, action, and transitions of $S$-Stealth Mining
Appendix C Proof of Theorem 2 and Algorithm 1